OpenStudy (help!!!!):

10. Melinda's lights went out. She has 3 pairs of red socks in her drawer, 2 pairs of black socks, and 5 pairs of white socks. What is the minimum number of pairs she must remove from the drawer to ensure that she has a pair of each color? A. 3 B. 5 C. 7 D. 9 E. 10 & how did you find the answer out?

OpenStudy (anonymous):

under worst case scenario first pair she draws will be white 2nd pair - white 3rd par - white 4th pair - white 5th pair - white 6th pair - red 7th pair - red 8th pair - red 9th pair - black When she draws the 9th time we notice that she has drawn a pair of each socks.

OpenStudy (help!!!!):

how did you know how to do it?

OpenStudy (help!!!!):

so if a similar problem appears the answer will be less than 1 from the total?

OpenStudy (anonymous):

worst case would give the highest number first... find the worst case by sorting max --> min for this question is 5, 3, 2 So, its 5*2+3*2+one pair of black = 10+6+2=18. 18/2 =9 pair

OpenStudy (help!!!!):

why mutiply by 2?

OpenStudy (anonymous):

2 socks makes 1 pair of socks

OpenStudy (help!!!!):

kk thanks